Of the appliance repair shops listed in the phone book, 80% are competent and 20% are...

  1. Of the appliance repair shops listed in the phone book,
    80% are competent and 20% are not. A competent shop can repair an appliance correctly 95% of the time; an incompetent shop can repair an appliance correctly 60% of the time.
    Let C = { shop is competent }
          R = { shop can repair appliance correctly } and find the probability that:

    (a) the shop is incompetent, given that the appliance was repaired correctly
    (b) the shop is incompetent, given that the appliance was not repaired correctly
    Hint: Draw a probability tree and use Bayes' formula:

2 .Suppose P(E) = 0.2, P(F) = 0.4, and P(E | F) = 0.3. Find the following probabilities. (a) PE'| F=
(b) PE | F'=
Answer #1

Solutions 11 P (C) = 0.8 P =0.95 P ( A ) = 0.60 P(R) = P(E). P() + P 1 . P() 0.8 +0.64 0.2 = = 0.95% 0.88 P(R) P(R) = 0.6x0.21 - 0.3 - 9) O PESO 1-P19) = 1-0.3 = 1017] b) PI & ) = P = TP (FC) PIEOFS = P1F) - P(EN F) = 0.4-0.3x0.4 1-0.4 - - 10.4667